OpenStudy (lgbasallote):

Two vectors \(\vec A\) and \(\vec B\) are given by \(\vec A = 5\hat i + 6\hat j + 7\hat k\) and \(\vec B = 3\hat i- 8\hat j + 2\hat k\). If these two vectors are drawn starting at the same point, what is the angle between them?

OpenStudy (anonymous):

use dot product formula...

OpenStudy (anonymous):

A . B= A B Cos(theta)

OpenStudy (lgbasallote):

can't i just use something simple like \[\tan^{-1} \left (\frac yx\right)\]

OpenStudy (anonymous):

what's y and x here?

OpenStudy (lgbasallote):

i have no idea....well i also tried getting the value of vector A and B...

OpenStudy (lgbasallote):

can i find out the answer using those info?

OpenStudy (anonymous):

actually if a vector R = x i + y j then theta is the the angle make by vector with x-axis

OpenStudy (anonymous):

which u wrote theta = arc tan(y/x)

OpenStudy (anonymous):

yeah to find A use the dispacement formula. thats root(i²+j²+k²)

OpenStudy (lgbasallote):

yes...so if i have A and B...what can i do next?

OpenStudy (lgbasallote):

im pretty sure it's not \[\tan^{-1} \left (\frac BA \right)\]

OpenStudy (anonymous):

A . B= A B Cos(theta)...use this...LHS is the dot product of vector A and B

OpenStudy (anonymous):

and On RHS ... A and B are the magnitudes of vectors A and B

OpenStudy (lgbasallote):

A and B are the values of A and B i found?

OpenStudy (lgbasallote):

the ones without the i j and k?

OpenStudy (anonymous):

lhs = 5.3+6.(-8)+7.2

OpenStudy (lgbasallote):

uhh where did these numbers come from?

OpenStudy (anonymous):

then use the cosine law